OpenStudy (pinkmilk234567):

4. How many solutions does the system of equations have? y = 4x + 3 and y = 4x + 4 (1 point) one two infinitely many none

OpenStudy (pinkmilk234567):

please help

OpenStudy (thecalchater):

can you simplify them?

OpenStudy (pinkmilk234567):

um i think

563blackghost (563blackghost):

These to equation have the same slope since they are in slope-intercept form which means that they are both parallel also becuase they have different y-intercepts....since they are parallel that would mean that they have no solution. http://www.montereyinstitute.org/courses/DevelopmentalMath/TEXTGROUP-1-19_RESOURCE/U14_L1_T1_text_container.html

OpenStudy (wolf1728):

Also since y = 4x + 3 and y = 4x + 4 That means that 4x +3 = 4x +4 So 4x - 4x = 1 Can you solve that?

OpenStudy (wolf1728):

Pinkmilk234567 can you solve the equation I just printed?

OpenStudy (pinkmilk234567):

sure

OpenStudy (wolf1728):

Okay please solve 4x - 4x = 1

OpenStudy (wolf1728):

4x - 4x = what?

OpenStudy (pinkmilk234567):

-16

OpenStudy (wolf1728):

No - think again

OpenStudy (pinkmilk234567):

16

OpenStudy (wolf1728):

No it is zero

OpenStudy (pinkmilk234567):

oh

OpenStudy (wolf1728):

So going back to our equation we get 0 = 1 which NEVER happens right?

OpenStudy (pinkmilk234567):

right

OpenStudy (wolf1728):

So the 2 equations can be reduced to 4x - 4x = 1 and 0 =1 So how many solutions does that give us?

OpenStudy (pinkmilk234567):

1 ?

OpenStudy (wolf1728):

No it has no solutions and that is your answer.

OpenStudy (pinkmilk234567):

oh okay can you help me with another

OpenStudy (wolf1728):

Well I have to go Perhaps someone else can help :-)

OpenStudy (pinkmilk234567):

okay well thank you for helping me ":)

OpenStudy (thecalchater):

still need help?